# Approximate Degrees of Multisymmetric Properties with Application to Quantum Claw Detection

### v1

@@ 4 Degree of Claw Property
This section provides a lower bound for the claw problem with a small range using Theorem 1. For this, we need a lower bound on the minimum degree of a polynomial that approximates the claw property for some common range size of input functions. To prove the following theorem, we use the Zhang’s reduction [32] from the collision problem for an input function that is either one-to-one or two-to-one.
@@
For any constant $20<\epsilon<1/$ 2, the minimum degree of a polynomial that $\epsilon$-approximates the claw property for function pairs in $[FG]^{F}\times[FG]^{G}$ is $\Omega\left((FG)^{1/3}\right)$.
@@
For disjoint subsets $S,T\subset[M]$, let $P^{\prime}_{ST}$ be a polynomial of degree $d$ (in variables $x_{ij}$ for $i\in S\sqcup T$ and $j\in[M]$) that $\epsilon$-approximates the claw property on function pairs $[M]^{S}\times[M]^{T}$. For every function $h$, let $P^{\prime}_{ST}(h)$ denote the value of $P^{\prime}_{ST}$ at the assignment to $x_{ij}$ corresponding to $h|_{S}$ and $h|_{T}$. Then, define a polynomial $P$ as the average of $P^{\prime}_{ST}$ over uniformly random disjoint subsets $S,T\subset[M]$ of size $F$ and $G$, respectively:
@@
If we set $M=FG$ for $G\geq F\geq 400$ and $\epsilon=0.1$, then it follows from Sec. missing  4 that $P(h)>0.2\cdot(1-0.1)=0.18$. Therefore, $P$ separates the one-to-one and two-to-one cases with a constant gap. By scaling and shifting $P$, we have a polynomial that $O(1)$-approximates the collision property. For instance, $\tilde{P}:=(25P+18)/43$ is sufficient and its approximation error is at most $41/86$. Therefore, $\deg(\tilde{P})=\Omega(M^{1/3})=\Omega((FG)^{1/3})$. Since $d\geq\deg(P)=\deg(\tilde{P})$, the minimum degree $d$ of a polynomial that 0.1-approximates the claw property for function pairs $[FG]^{F}\times[FG]^{G}$ is $\Omega\left((FG)^{1/3}\right)$. By Sec. missing  2.2, the theorem holds. $\square$
@@
Proof.  To choose random disjoint subsets $S$ and $T$, we first choose a random subset $S$ of $[M]$ and then choose a random subset $T$ of $[M]\setminus S$. Observe that, if $h|_{S}$ is injective and, in addition, $h^{-1}(h(S))\setminus S$ intersects $T$, then a claw exists. Thus, we obtain
@@
where the first inequality follows from $1+x\geq\exp(\frac{x}{1+x})$, and the last inequality follows from $1004F\leq F^{2}/$ 100 for $F\geq 400$. For the last formula to be at least $1-\delta$, we have
@@
As a corollary of Theorem 12, we obtain the following theorem.
@@
Let $F,G\in\mathbb{N}$ be such that $F\leq G$. For any constant $20<\epsilon<1/$ 2 and any $M\geq F+G$, the minimum degree of a polynomial that $\epsilon$-approximates the claw property for function pairs in $[M]^{[F]}\times[M]^{[G]}$ is $\Theta(\sqrt{G}+(FG)^{1/3})$, where the lower bound $\Omega(\sqrt{G})$ holds for all $M\geq 2$.
Proof.  Let $\Phi$ be the claw property for functions in $[FG]^{[F]}\times[FG]^{[G]}$. By Theorem 16, we have $\widetilde{\deg}_{\epsilon}(\Phi)=\Omega((FG)^{1/3})$ when $M=FG$. By applying Theorem 12, we have $\widetilde{\deg}_{\epsilon}(\Phi^{\prime})=\Omega((FG)^{1/3})$, where $\Phi^{\prime}$ is the restriction of $\Phi$ to the domain $[M^{\prime}]^{[F]}\times[M^{\prime}]^{[G]}$ for any $M^{\prime}$ with $F+G\leq M^{\prime}<FG$. This proves $\Omega((FG)^{1/3})$ lower bound in the statement of the theorem.
@@
Since there is an $O\left(\sqrt{G}+(FG)^{1/3}\right)$-query quantum algorithm that detects a claw (and finds it if it exists) [30], the polynomial method (Sec. missing  2.2) implies that the lower bound $\Omega\left(\sqrt{G}+(FG)^{1/3}\right)$ is tight for $M\geq F+G$. $\square$
The proof of Theorem 18 implies the optimal quantum query complexity as follows.
@@
Let $F,G\in\mathbb{N}$ be such that $F\leq G$. For every $M\geq F+G$, the (query-)optimal quantum algorithm that computes a claw (or detects the existence of a claw) with a constant error probability requires $\Theta(\sqrt{G}+(FG)^{1/3})$ queries for a given pair of functions in $[M]^{[F]}\times[M]^{[G]}$. More precisely, the optimal quantum query complexity is $\Theta((FG)^{1/3})$ for every $M\geq F+G$ if $F\leq G\leq F^{2}$ and $\Theta(\sqrt{G})$ for every $M\geq 2$ if $G>F^{2}$.
@@ 5 Claw Problem with Much Smaller Range
This section proves Theorem 5. The proof generalizes the idea used in the case of the same domain size [7]: We first reduce the composition of a certain problem (called $\mathsf{pSearch}$) and the claw problem for a larger size of the function range to the claw problem in question, and then use the composition theorem of the adversary method.
@@
By employing Theorem 19 and Sec. missing  5, we obtain the following theorem.
@@
Let $F,G,M\in\mathbb{N}$ be such that $F\leq G\leq F^{2}$ and $M<F+G$ [^2] The condition of $F\leq G\leq F^{2}$ and $M<F+G$ means the parameter range that Theorem 19 does not cover. Then, for every $M\in[2,F+G-1]$, the quantum query complexity of detecting the existence of a claw for a given function pair $(f,g)$ in $[M]^{[F]}\times[M]^{[G]}$ is lower-bounded by
@@
The input to $\mathcal{P}$ is a pair of functions $f:=(f_{1},\dots,f_{F})\in([M]\cup\{\ast\})^{[F]}$ and $g:=(g_{1},\dots,g_{G})\in([M]\cup\{\ast\})^{[G]}$ (as the sequence representation of functions; see Sec. missing  2.1) for $M\geq 2$ with the following promise, where $f_{i}$ denotes $f(i)$ for $i\in[F]$ and $g_{k}$ denotes $g(k)$ for $k\in[G]$:
@@
where $\mathsf{claw}_{(a,b)\to c}$is the decision problem of detecting the existence of a claw for given functions $(\phi,\psi)\in[c]^{[a]}\times[c]^{[b]}$. Note that, since each query access to $f^{\prime}\ (g^{\prime})$ can be simulated by a single query access to $f\ (\text{respectively},\ g)$, the modification to input functions needs no extra queries. Therefore, the quantum query complexity of $\mathsf{claw}_{(F,G)\to M+2}$ is lower-bounded by the quantum query complexity of $\mathsf{claw}_{(k,kG/F)\to M}\circ\mathsf{pSearch}_{F/k\to M}$.
@@
By applying Theorem 20 and Sec. missing  5 with Eq. (missing) 4 and Eq. (missing) 5, we have
@@
where the first equality follows from Eq. (missing) 6 and the last equality follows from $F^{\prime}\geq F/2$ and $G^{\prime}\geq G/2$ as shown in Sec. missing  5.
Next, we will remove the assumption that $M$ is of the form $M=k+kG^{\prime}/F^{\prime}$ and show that the lower bound holds for every integer $M$ such that $2\leq M<F+G$. For this, let us fix $F$ and $G$ arbitrarily such that $F\leq G\leq F^{2}$. For every integer $M$ such that $M\in\left[1+\left\lfloor G/F\right\rfloor,F+G\right]$, there exists $k\in[1,F]$ such that $M_{k}\leq M\leq M_{k+1}$ by Sec. missing  5, where $M_{k}$ is defined in the claim. Since $\mathsf{claw}_{\left(F,G\right)\rightarrow M_{k}+2}$ is reducible to $\mathsf{claw}_{\left(F,G\right)\rightarrow M+2}$ since $[M_{k}+2]\subseteq[M+2]$, we have
@@
where the last equality follows from $M/M_{k}<M_{k+1}/M_{k}\leq 8$ by Sec. missing  5. By replacing $M+2$ with $M$, we have
@@
for $3+\left\lfloor G/F\right\rfloor\leq M\leq F+G$. Since $\Omega\left(\sqrt{G}+F^{1/3}G^{1/6}M^{1/6}\right)=\Omega\left(\sqrt{G}\right)$ for every positive $M<3+\left\lfloor G/F\right\rfloor$ and $Q_{1/3}\left(\mathsf{claw}_{\left(F,G\right)\rightarrow M}\right)=\Omega\left(\sqrt{G}\right)$ [14] for every $M\geq 2$ follows from the unstructured search, it holds that $Q_{1/3}\left(\mathsf{claw}_{\left(F,G\right)\rightarrow M}\right)=\Omega\left(\sqrt{G}+F^{1/3}G^{1/6}M^{1/6}\right)$ for $2\leq M\leq F+G$. More precisely, Since $\sqrt{G}>F^{1/3}G^{1/6}M^{1/6}$ if and only if $\left(G/F\right)^{2}>M$, we have [^3] The case $2\leq M<(G/F)^{2}$ should be ignored when $G/F\leq\sqrt{2}$.
@@
We next prove the first and the last items. Since $F_{k}^{\prime}\leq F$ and $G_{k}^{\prime}\leq G$ by the definition, the first item follows by proving the lower bounds on $F^{\prime}_{k}$ and $G^{\prime}_{k}$. By the definition of $F_{k}^{\prime}$, we have $F-F_{k}^{\prime}<\left\lfloor F/k\right\rfloor$ and $F_{k}^{\prime}\geq\left\lfloor F/k\right\rfloor$. Hence,
@@ Appendix A Claim of Polynomial Method
where $+$ is the addition modulo $M$, or alternatively,
@@
where $\oplus$ is bitwise XOR in binary expression. At any step, the the state of the algorithm is of the form
@@
Proof.  For $q=0$, since $\alpha_{s,i,b,z}$ is independent of the oracles, the statement holds. Suppose that the claim holds for $q-1$ and let $\sum_{s,i,b,z}{\alpha_{s,i,b,z}\left|s,i\right\rangle\left|b\right\rangle\left|z\right\rangle}$ be the state after applying $U_{q-1}$, where each $\alpha_{s,i,b,z}$ is a polynomial over $x_{ij},y_{ij}$ of degree at most $q-1$. Assume that the $q$th $O_{f/g}$ maps this state to $\sum_{s,i,b,z}{\beta_{s,i,b,z}\left|s,i\right\rangle\left|b\right\rangle\left|z\right\rangle}$. Observe that $O_{f/g}$ maps each basis state $|0,i\rangle|b\rangle|z\rangle$ to $|0,i\rangle|b+f(i)\rangle|z\rangle$ (likewise $|1,k\rangle|b\rangle|z\rangle$) to $|0,k\rangle|b+g(k)\rangle|z\rangle$), and that $|0,i\rangle|b-j+f(i)\rangle|z\rangle$ is exactly $|0,i\rangle|b\rangle|z\rangle$ if $x_{ij}=1$ (and likewise $|1,k\rangle|b-j+g(k)\rangle|z\rangle$ is exactly $|1,k\rangle|b\rangle|z\rangle$ if $y_{kj}=1$). Thus, it holds that
@@
Lemma 26 (restatement of Sec. missing  2.2)
Let $F,G,M$ be in $\mathbb{N}$, and let $\Phi$ be a Boolean function over the set of pairs of functions in $[M]^{[F]}\times[M]^{[G]}$. For given oracles $O_{f}$ and $O_{g}$ for functions $f\in[M]^{[F]}$ and $g\in[M]^{[G]}$, suppose that there exists a quantum algorithm that computes $\Phi(f,g)$ with error $\epsilon$ using $q$ queries. Then, there is a polynomial $P$ of degree at most $2q$ in variables $x_{11},\dots,x_{FM},y_{11},\dots,y_{GM}$ such that $P$ $\epsilon$-approximates $\Phi$, where $x_{ij}$ and $y_{kj}$ are the variables defined for $f$ and $g$, respectively, in Sec. missing  2.2.
Proof.  Assume without loss of generality that the quantum algorithm measures the last single-qubit register in the final state $\sum_{s,i,b,z}{\alpha_{s,i,b,z}\left|s,i\right\rangle\left|b\right\rangle\left|z\right\rangle}$ in the computational basis and outputs the outcome as the value of $\Phi$. Let $A$ be the set of tuples $(s,i,b,z)$ with the last bit of $z$ being $\Phi(f,g)$. Then, the probability that the algorithm outputs $\Phi(f,g)$ is
@@
It follows from Sec. missing  2.2 that $p_{\Phi}$ is represented as a polynomial of degree at most $2q$ in $x_{11},\dots,x_{FM},y_{11},\dots,y_{GM}$. Since the error probability of the algorithm is at most $\epsilon$, we have $\left\lvert p_{\Phi}(f,g)-\Phi(f,g)\right\rvert\leq\epsilon$ for all $f,g$. This implies that there is a polynomial $P$ of degree at most $2q$ in variables $x_{11},\dots,x_{FM},y_{11},\dots,y_{GM}$ such that $P$ $\epsilon$-approximates $\Phi$. $\square$
Lemma 27 ([15]; restatement of Sec. missing  2.2 )
For a partial Boolean function $\Phi$ and $20<\epsilon<\delta<1/$ 2, it holds that
@@
Intuitively, this is the probability of seeing at least $\ell/2$ heads among $\ell$ coin flips, if $p$ is the probability of seeing the heads when flipping a single coin. By Chernoff’s bound, the approximation error of $A(p)$ is $\exp(-2\ell(1/2-\delta)^{2})$. For this to be $\epsilon$, we have

### v2

@@ 4 Degree of Claw Property
This section provides a lower bound for the claw problem with a small range using Theorem 1. For this, we need a lower bound on the minimum degree of a polynomial that approximates the claw property for some common range size of input functions. To prove the following theorem, we use Zhang’s reduction [33] from the collision problem for an input function that is either one-to-one or two-to-one.
@@
For any constant $0<\epsilon<1/2$, the minimum degree of a polynomial that $\epsilon$-approximates the claw property for function pairs in $FG]^{F}\times[FG]^{G}$ is $\Omega\left((FG)^{1/3}\right)$.
@@
For disjoint subsets $S,T\subset[M]$, let $P^{\prime}_{ST}$ be a polynomial of degree $d$ (in variables $x_{ij}$ for $i\in S\sqcup T$ and $j\in[M]$) that $\epsilon$-approximates the claw property on function pairs $M]^{S}\times[M]^{T}$. For every function $h$, let $P^{\prime}_{ST}(h)$ denote the value of $P^{\prime}_{ST}$ at the assignment to $x_{ij}$ corresponding to $h|_{S}$ and $h|_{T}$. Then, define a polynomial $P$ as the average of $P^{\prime}_{ST}$ over uniformly random disjoint subsets $S,T\subset[M]$ of size $F$ and $G$, respectively:
@@
[…]
@@
Proof. To choose random disjoint subsets $S$ and $T$, we first choose a random subset $S$ of $[M]$ and then choose a random subset $T$ of $M]\setminus S$. Observe that, if $h|_{S}$ is injective and, in addition, $h^{-1}(h(S))\setminus S$ intersects $T$, then a claw exists. Thus, we obtain
@@
where the first inequality follows from $1+x\geq\exp(\frac{x}{1+x})$, and the last inequality follows from $4F\leq F^{2}/100$ for $F\geq 400$. For the last formula to be at least $1-\delta$, we have
@@
We obtain the following theorem as a corollary of Theorem 12.
@@
Let $F,G\in\mathbb{N}$ be such that $F\leq G$. For any constant $0<\epsilon<1/2$ and any $M\geq F+G$, the minimum degree of a polynomial that $\epsilon$-approximates the claw property for function pairs in $M]^{[F]}\times[M]^{[G]}$ is $\Theta(\sqrt{G}+(FG)^{1/3})$, where the lower bound $\Omega(\sqrt{G})$ holds for all $M\geq 2$.
Proof. Let $\Phi$ be the claw property for functions in $FG]^{[F]}\times[FG]^{[G]}$. By Theorem 16, we have $\widetilde{\deg}_{\epsilon}(\Phi)=\Omega((FG)^{1/3})$ when $M=FG$. By applying Theorem 12, we have $\widetilde{\deg}_{\epsilon}(\Phi^{\prime})=\Omega((FG)^{1/3})$, where $\Phi^{\prime}$ is the restriction of $\Phi$ to the domain $M^{\prime}]^{[F]}\times[M^{\prime}]^{[G]}$ for any $M^{\prime}$ with $F+G\leq M^{\prime}<FG$. This proves a $\Omega((FG)^{1/3})$ lower bound in the statement of the theorem.
@@
Since there is an $O\left(\sqrt{G}+(FG)^{1/3}\right)$-query quantum algorithm that detects a claw (and finds it if it exists) [31], the polynomial method (Sec.   2.2) implies that the lower bound $\Omega\left(\sqrt{G}+(FG)^{1/3}\right)$ is tight for $M\geq F+G$. $\square$
The proof of Theorem 18 implies the following optimal quantum query complexity.
@@
Let $F,G\in\mathbb{N}$ be such that $F\leq G$. For every $M\geq F+G$, the (query-)optimal quantum algorithm that computes a claw (or detects the existence of a claw) with a constant error probability requires $\Theta(\sqrt{G}+(FG)^{1/3})$ queries for a given pair of functions in $M]^{[F]}\times[M]^{[G]}$. More precisely, the optimal quantum query complexity is $\Theta((FG)^{1/3})$ for every $M\geq F+G$ if $F\leq G\leq F^{2}$ and $\Theta(\sqrt{G})$ for every $M\geq 2$ if $G>F^{2}$.
@@ 5 Claw Problem with Much Smaller Range
This section proves Theorem 5. The proof generalizes the idea used in the case of the same domain size [7]: We first reduce the composition of a certain problem (called $\mathsf{pSearch}$) and the claw problem for a larger size of the function range, to the claw problem in question, and then use the composition theorem of the adversary method.
@@
By employing Theorem 19 and Sec.   5, we obtain the following theorem.
@@
Let $F,G,M\in\mathbb{N}$ be such that $F\leq G\leq F^{2}$ and M<F+GM<F+G. [^2] The condition of $F\leq G\leq F^{2}$ and $M<F+G$ means the parameter range that Theorem 19 does not cover. Then, for every $M\in[2,F+G-1]$, the quantum query complexity of detecting the existence of a claw for a given function pair $(f,g)$ in $M]^{[F]}\times[M]^{[G]}$ is lower-bounded by
@@
The input to $\mathcal{P}$ is a pair of functions $f:=(f_{1},\dots,f_{F})\in([M]\cup\{\ast\})^{[F]}$ and $g:=(g_{1},\dots,g_{G})\in([M]\cup\{\ast\})^{[G]}$ (as the sequence representation of functions; see Sec.   2.1) for $M\geq 2$ with the following promise, where $f_{i}$ denotes $f(i)$ for $i\in[F]$ and $g_{k}$ denotes $g(k)$ for $k\in[G]$:
@@
where $\mathsf{claw}_{(a,b)\to c}$is the decision problem of detecting the existence of a claw for given functions $(\phi,\psi)\in[c]^{[a]}\times[c]^{[b]}$. Note that, since each query access to $f^{\prime}\ (g^{\prime})$ can be simulated by a single query access to f(respectively,g)f\ ($\text${respectively},\ g), the modification to input functions needs no extra queries. Therefore, the quantum query complexity of $\mathsf{claw}_{(F,G)\to M+2}$ is lower-bounded by the quantum query complexity of $\mathsf{claw}_{(k,kG/F)\to M}\circ\mathsf{pSearch}_{F/k\to M}$.
@@
By applying Theorem 20 and Sec.   5 with Eq. () 4 and Eq. () 5, we have
@@
where the first equality follows from Eq. () 6 and the last equality follows from $F^{\prime}\geq F/2$ and $G^{\prime}\geq G/2$ as shown in Sec.   5.
Next, we will remove the assumption that $M$ is of the form $M=k+kG^{\prime}/F^{\prime}$ and show that the lower bound holds for every integer $M$ such that $2\leq M<F+G$. For this, let us arbitrarily fix $F$ and $G$ such that $F\leq G\leq F^{2}$. For every integer $M$ such that $M\in\left[1+\left\lfloor G/F\right\rfloor,F+G\right]$, there exists $k\in[1,F]$ such that $M_{k}\leq M\leq M_{k+1}$ by Sec.   5, where $M_{k}$ is defined in the claim. Since $\mathsf{claw}_{\left(F,G\right)\rightarrow M_{k}+2}$ is reducible to $\mathsf{claw}_{\left(F,G\right)\rightarrow M+2}$ since $M_{k}+2]\subseteq[M+2]$, we have
@@
where the last equality follows from $M/M_{k}<M_{k+1}/M_{k}\leq 8$ by Sec.   5. By replacing $M+2$ with $M$, we have
@@
for $3+\left\lfloor G/F\right\rfloor\leq M\leq F+G$. Since it holds that $\Omega\left(\sqrt{G}+F^{1/3}G^{1/6}M^{1/6}\right)=\Omega\left(\sqrt{G}\right)$ for every positive $M<3+\left\lfloor G/F\right\rfloor$, and it follows from the unstructured search that $Q_{1/3}\left(\mathsf{claw}_{\left(F,G\right)\rightarrow M}\right)=\Omega\left(\sqrt{G}\right)$ [14] for every $M\geq 2$, it holds that $Q_{1/3}\left(\mathsf{claw}_{\left(F,G\right)\rightarrow M}\right)=\Omega\left(\sqrt{G}+F^{1/3}G^{1/6}M^{1/6}\right)$ for $2\leq M\leq F+G$. More precisely, Since $\sqrt{G}>F^{1/3}G^{1/6}M^{1/6}$ if and only if $\left(G/F\right)^{2}>M$, we have [^3] The case $2\leq M<(G/F)^{2}$ should be ignored when $G/F\leq\sqrt{2}$.
@@
We next prove the first and the last items. Since $F_{k}^{\prime}\leq F$ and $G_{k}^{\prime}\leq G$ by the definition, the first item follows from the lower bounds on $F^{\prime}_{k}$ and $G^{\prime}_{k}$. By the definition of $F_{k}^{\prime}$, we have $F-F_{k}^{\prime}<\left\lfloor F/k\right\rfloor$ and $F_{k}^{\prime}\geq\left\lfloor F/k\right\rfloor$. Hence,
@@ Appendix A Claim of Polynomial Method
where ++ is the addition modulo $M$, or alternatively,
@@
where $\oplus$ is bitwise XOR in binary expression. At any step, the state of the algorithm is of the form
@@
Proof. For $q=0$, since $\alpha_{s,i,b,z}$ is independent of the oracles, the statement holds. Suppose that the claim holds for q1q-1 and let $\sum_{s,i,b,z}{\alpha_{s,i,b,z}\left|s,i\right\rangle\left|b\right\rangle\left|z\right\rangle}$ be the state after applying $U_{q-1}$, where each $\alpha_{s,i,b,z}$ is a polynomial over $x_{ij},y_{ij}$ of degree at most q1q-1. Assume that the $q$th $O_{f/g}$ maps this state to $\sum_{s,i,b,z}{\beta_{s,i,b,z}\left|s,i\right\rangle\left|b\right\rangle\left|z\right\rangle}$. Observe that $O_{f/g}$ maps each basis state $|0,i\rangle|b\rangle|z\rangle$ to $|0,i\rangle|b+f(i)\rangle|z\rangle$ (likewise $|1,k\rangle|b\rangle|z\rangle$) to $|0,k\rangle|b+g(k)\rangle|z\rangle$), and that $|0,i\rangle|b-j+f(i)\rangle|z\rangle$ is exactly $|0,i\rangle|b\rangle|z\rangle$ if $x_{ij}=1$ (and likewise $|1,k\rangle|b-j+g(k)\rangle|z\rangle$ is exactly $|1,k\rangle|b\rangle|z\rangle$ if $y_{kj}=1$). Thus, it holds that
@@
Lemma 26 (restatement of Sec.   2.2)
Let $F,G,M$ be in $\mathbb{N}$, and let $\Phi$ be a Boolean function over the set of pairs of functions in $M]^{[F]}\times[M]^{[G]}$. For given oracles $O_{f}$ and $O_{g}$ for functions $f\in[M]^{[F]}$ and $g\in[M]^{[G]}$, suppose that there exists a quantum algorithm that computes $\Phi(f,g)$ with error $\epsilon$ using $q$ queries. Then, there is a polynomial $P$ of degree at most $2q$ in variables $x_{11},\dots,x_{FM},y_{11},\dots,y_{GM}$ such that $P$ $\epsilon$-approximates $\Phi$, where $x_{ij}$ and $y_{kj}$ are the variables defined for $f$ and $g$, respectively, in Sec.   2.2.
Proof. Assume, without loss of generality, that the quantum algorithm measures the last single-qubit register in the final state $\sum_{s,i,b,z}{\alpha_{s,i,b,z}\left|s,i\right\rangle\left|b\right\rangle\left|z\right\rangle}$ in the computational basis and outputs the outcome as the value of $\Phi$. Let $A$ be the set of tuples $(s,i,b,z)$ with the last bit of $z$ being $\Phi(f,g)$. Then, the probability that the algorithm outputs $\Phi(f,g)$ is
@@
It follows from Sec.   2.2 that $p_{\Phi}$ is represented as a polynomial of degree at most $2q$ in $x_{11},\dots,x_{FM},y_{11},\dots,y_{GM}$. Since the error probability of the algorithm is at most $\epsilon$, we have $\left\lvert p_{\Phi}(f,g)-\Phi(f,g)\right\rvert\leq\epsilon$ for all $f,g$. This implies that there is a polynomial $P$ of degree at most $2q$ in variables $x_{11},\dots,x_{FM},y_{11},\dots,y_{GM}$ such that $P$ $\epsilon$-approximates $\Phi$. $\square$
Lemma 27 ([16]; restatement of Sec.   2.2 )
For a partial Boolean function $\Phi$ and $0<\epsilon<\delta<1/2$, it holds that
@@
Intuitively, this is the probability of seeing at least $\ell/2$ heads among $\ell$ coin flips, if $p$ is the probability of seeing heads when flipping a single coin. By Chernoff’s bound, the approximation error of $A(p)$ is $\exp(-2\ell(1/2-\delta)^{2})$. For this to be $\epsilon$, we have